# Contrastive Learning and Adversarial Disentanglement for Privacy-Preserving Task-Oriented Semantic Communications

### v1

@@ VI-D2 Colored FashionMNIST Results
In Table IV CLAD outperforms DeepJSCC with a 2.46% improvement and VIB at $\beta=0.001$ with 2.44% increase in accuracy. The IRI and adversarial accuracy results for Colored FashionMNIST reflect similar trends as observed in the Colored MNIST dataset. CLAD achieves a considerably lower IRI and lower adversarial accuracy compared to DeepJSCC and VIB at $\beta=0.001$, further confirming that CLAD is better at disentangling task-irrelevant information and preserving privacy. Fig. 6(b), visually illustrates the varying levels of informativeness and minimality in the encoded representations produced by each method, with respect to the input. In this image CLAD effectively removes most style features such as color, texture, and fine details, focusing solely on the basic shapes necessary for classification. For example, the CLAD reconstructions of the shoe, trousers, and shirt are reduced to simple grayscale silhouettes, discarding the vibrant colors and textures seen in the original and DeepJSCC versions. DeepJSCC retains style features such as color and subtle details such as the texture of fabric, making the reconstructions more complex but less focused on task-relevant information. VIB also retains some color and texture but introduces significant blurring, partially losing both class-relevant and style features. In contrast, CLAD’s grayscale, simplified images prioritize the task-relevant information necessary for downstream classification.
Once again, from Table IV, VIB achieves smaller IRI values at smaller $\beta$ settings but suffers a drop in accuracy, and the IRI continues to decrease as $\beta$ increases at the expense of even lower accuracy. This demonstrates that while VIB can be tuned to prioritize privacy and improve the minimality of the feature vector, it often struggles to maintain high classification performance.

### v2

@@ VI-D2 Colored FashionMNIST Results
In Table VI CLAD outperforms DeepJSCC with a 2.46% improvement and VIB at $\beta=0.001$ with 2.44% increase in accuracy. The IRI and adversarial accuracy results for Colored FashionMNIST reflect similar trends as observed in the Colored MNIST dataset. CLAD achieves a considerably lower IRI and lower adversarial accuracy compared to DeepJSCC and VIB at $\beta=0.001$, further confirming that CLAD is better at disentangling task-irrelevant information and maintaining privacy awareness. Fig. LABEL:fig:fmnist_reconstructions, visually illustrates the varying levels of informativeness and minimality in the encoded representations produced by each method, with respect to the input. In this image CLAD effectively removes most style features such as color, texture, and fine details, focusing solely on the basic shapes necessary for classification. For example, the CLAD reconstructions of the shoe, trousers, and shirt are reduced to simple grayscale silhouettes, discarding the vibrant colors and textures seen in the original and DeepJSCC versions. DeepJSCC retains style features such as color and subtle details such as the texture of fabric, making the reconstructions more complex but less focused on task-relevant information. VIB also retains some color and texture but introduces significant blurring, partially losing both class-relevant and style features. In contrast, CLAD’s grayscale, simplified images prioritize the task-relevant information necessary for downstream classification.
Once again, from Table VI, VIB achieves smaller IRI values at smaller $\beta$ settings but suffers a drop in accuracy, and the IRI continues to decrease as $\beta$ increases at the expense of even lower accuracy. This demonstrates that while VIB can be tuned to prioritize privacy and improve the minimality of the feature vector, it often struggles to maintain high classification performance.
VI-D3 CIFAR-10 Results
For the more complex CIFAR-10 dataset, Fig. 4(c) illustrates that CLAD continues to outperform baseline methods across all SNR levels. The performance gap becomes particularly notable at higher SNRs, where CLAD achieves up to 2% improvement in classification accuracy over DeepJSCC and VIB.
As shown in Table VII, CLAD delivers the highest classification accuracy of 92.33% at 12 dB, outperforming DeepJSCC by 1.13% and VIB (with $\beta=0.001$) by 0.84%. Moreover, CLAD achieves a significantly lower IRI compared to DeepJSCC, demonstrating its ability to retain task-relevant features while discarding redundant information, even in a dataset with rich visual complexity and no explicit sensitive attributes. These results reinforce CLAD’s robustness in encoding informative yet compact representations for more challenging computer vision tasks, without compromising classification performance.